OpenStudy (anonymous):

What is the solution to the equation? 4x – 6 = –3x + 8 A. 14 B. 7/2 C. 2/7 D. 2

OpenStudy (anonymous):

4x - 6=-3x + 8 First, we will add 3x to both sides 4x - 6 + 3x = -3x + 8 + 3x or 7x - 6= 8 Next we will add 6 to both sides, giving us 7x = 14 Finally, we will divide both sides by 7, giving us x=2, so the answer is D.

OpenStudy (anonymous):

Thanks. I understand it now.

OpenStudy (anonymous):

You're welcome. Glad to be of assistance.